To save for retirement, you deposit $8,592.06 at the end of every 6 months. If you make these payments for the next 44 years while you work, and the interest rate is 12.03%, how much will you have in your retirement account when you go to retire? That is, immeidately after the last payment is made

Answer =

Cash Flow =

r =

n =
Answer #1

Balance in retirement account when go to retire will be $ 2,42,45,686.18

Future value of cash flow = Semi annual cash flow * Future value of annuity of 1
= $             8,592.06 * 2821.871144
= $ 2,42,45,686.18
Working:
Future value of annuity of 1 = (((1+i)^n)-1)/i Where,
= (((1+0.06015)^88)-1)/0.06015 i = 12.03%/2 = 0.06015
= 2821.871144 n = 44*2 = 88